1.      Suppose A={4,3,6,7,1,9}, B={5,6,8,4} and C={5,8,4}. Find:

(a)   A ∪ B

(e) B-A

(h) B ∪ C

3. Suppose A={0,1} and B={1,2}. Find:

(a) (AxB) ∩(BxB)

(d) (A∩B)xA

(i) ℙ(AxB)


6. Sketch the sets X=[-1,3]x[0,2] and Y=[0,3]x[1,4] on the plane R2. On separate drawings, shade in the sets X∪Y, X∩Y, X-Y and Y-X.


10. Do you think the statement (R-Z)xN=(RxN)-(Z-N) is true, or false? Justify.


1
Expert's answer
2021-03-17T12:37:19-0400
  1. Suppose "A={4,3,6,7,1,9}," "B={5,6,8,4}" and "C={5,8,4}." Find:

(a)  "A \u222a B = {1,3,4,5,6,7,8,9}"

(e) "B-A = {5,8}"

(h) "B \u222a C = {5,6,8,4}"

3. Suppose "A={0,1}" and "B={1,2}" . Find:

(a) "(AxB) \u2229(BxB) = {(1,1),(1,2)}"

(d) "(A\u2229B)xA = {(1,0),(1,1)}"

(i) "ℙ(AxB) = \empty ,{(0,1)},{(0,2)},{(1,1)},{(1,2)},{(0,1),(0,2)},{(0,1),(1,1)},{(0,1),(1,2)},{(0,2),(1,1)},"

"{(0,2),(1,2)},{(1,1),(1,2)},{(0,2),(1,1),(1,2)},{(0,1),(1,1),(1,2)},{(0,1),(0,2),(1,2)},"

"{(0,1),(0,2),(1,1)}, {(0,1),(0,2),(1,1),(1,2)}"


6.



10.No because "(R-Z)xN = (RxN) - (ZxN)"

which means that "(RxN) - (ZxN)" is not equal to "(RxN)-(Z-N)"